OpenStudy (anonymous):

The perimeter of a triangle is 7a-11b. If two of its sides are 2a+b and a-9b what is the third side?

OpenStudy (anonymous):

Let the third side be A perimeter of triangle = sum of all its three sides i.e. 7a-11b = 2a+b + a-9b +A i.e. 7a-11b = 3a-8b +A i.e. 7a-11b -3a+8b =A i.e. 4a-3b =A Therefore third side of the given triangle= 4a-3b